OpenStudy (anonymous):

Which equation represents a line through (3, 5) that is perpendicular to y = 2x - 5.

OpenStudy (anonymous):

2y = -x + 13 2y - x = 13 2y = x + 13 2y + x = -13

jimthompson5910 (jim_thompson5910):

What is the slope of the line y = 2x - 5 ?

jimthompson5910 (jim_thompson5910):

Do you see what the slope of y = 2x - 5 is?